OpenStudy (anonymous):

how do you factor 3x^2+5x+2

OpenStudy (domebotnos):

3x^2 + 5x + 2 First multiply the constant by the co-efficient of the x^2 term. 3*2=6 Find the pair of factors that add up to 5, the middle terms co-efficient. 3*2=6 3+2=5 Then you pair them up like this: (x + 3)(x + 2) Then divide the co-efficient by 3 (x + 3)(x + 2) ___ ____ 3 3 3 goes into 3 evenly but not 2, so put the 3 with the x. (x + 1)(3x + 2) The answer is (x + 1)(3x + 2).